Medication order: Loading dose, mix magnesium sulfate 30 g in 1000 mL D5W. Infuse 6 g over 20 minutes, then maintain 1 g/hr. Maintain primary IV 1000 mL D5W at 60 mL/hr after magnesium sulfate loading dose.
Available: Magnesium sulfate 50% in 10 mL vials
Calculate the secondary rate for the loading dose.
___________
Answer #1

Step 1:

Order: Mix magnesium sulfate 30g in 1000 mL.

1000 mL = 30 g

Infusion order = 6 g over 20 minutes.

Therefore,

Amount = Ordered dose / Dose on hand X volume

= 6 / 30 X 1000

Amount = 200 mL

Hence 200 mL contains 6 g of magnesium sulfate to be infused over 20 minutes.

Step 2:

1 hour = 60 minutes.

Ordered time = 20 minutes

mL/hr = 60 / 20 X 200

mL/hr = 600

The infusion pump to be set 600 mL/hr to infuse 6 g of magnesium sulfate for over 20 minutes.

Step 3:

Ordered secondary loading dose = 1g/hr.

200 mL contains 6g

Therefore 1g = 200 / 6

1g = 33.3 mL/hr

or

Ordered dose = 1g

Available dose = 30g

Volume = 1000 mL

Amount = Ordered dose / Dose on hand X Volume

Amount = 1 / 30 X 1000

Amount = 33.3 mL /hr

The infusion should be set as 33.3 mL/hr for secondary loading Dose.